Suppose a firm producing coffee provides you with the following information:

Price = R100

Quantity = 150 units of coffee

Cost of production = R9 500


7.3.1. Calculate the total revenue of the firm. (1)


7.3.2. Is the firm currently making a profit or a loss? Explain. (2)


1
Expert's answer
2022-04-24T17:25:22-0400

The total revenue of the firm is:

"TR = 100\u00d7150 = R15,000."

Total profit is:

"TP = TR - TC = 15,000 - 9,500 = R5,500."

So, the firm is currently making a profit.